OpenStudy (anonymous):

A cylinder has a volume of 45 cubic inches. What is the Volume of a cone that the cone fits exactly inside of? Please help me!

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

1/3 of volume of cylinder.

OpenStudy (anonymous):

volume of cylinder= pi r^2h volume of cone=1/3 pi r^2h

OpenStudy (anonymous):

But how do I plug it in?

OpenStudy (anonymous):

\[volume ~of~cone=\frac{ 1 }{ 3 } \pi r^2 h=\frac{ 1 }{ 3 } volume ~of ~cylinder=\frac{ 1 }{ 3 } \times 45=?\]

OpenStudy (anonymous):

Im still lost..